# On total transitivity of graphs

### v1

@@ 3 Total Transitivity in split graphs
$|dom_{K}(S)|=2$
In this case, let $l_{min}=|dom_{K}(S)|$. By the definition of $dom_{K}(S)$, the vertices are in $V_{1}$. As we know from the Observation15, each $V_{i}$ must contain a vertex from $K$, so each $(p-1)-[q-\{(p-1)+l_{min}\}]=2p-q+l_{min}-2$ set from $\pi$ must contain at least one vertex from $S$. Therefore, $|S|\geq 2p-q+l_{min}-2$ $(b(i))$. Again, by Observation16, each $|V_{i}\cap S|\leq 2$. Let $\{s_{i_{1}},s_{i_{2}},\ldots,s_{i_{\beta}}\}\subseteq S$, where $\beta=2p-q+l_{min}-2$ such that $s_{i_{j}}\in V_{i_{j}}$. Also, note that $i_{j}\geq 2$. Since $\pi$ is a total transitive partition, $deg(s_{i_{j}})\geq i_{j}\geq j$ $(b(ii))$.
@@ 4 NP-completeness
In this section, we show that the Maximum Total Transitivity Decision Problem is NP-complete for a bipartite graph. Clearly, this problem is in NP. We prove the NP-hardness of this problem by showing a polynomial-time reduction from Proper $3$-Coloring Decision Problem in graphs to the MtotTDP. A proper $3$-colring of a graph $G=(V,E)$ is a function $g:V\rightarrow\{1,2,3\}$, such that for every edge $uv\in E$, $g(u)\not=g(v)$. The proper $3$-Coloring Decision Problem is defined as follows:
@@
Figure 3: Construction of $G^{\prime}$ from $G$ where $m=3$ and $T$ represents a $tcmbt_{3}$
The construction is as follows. Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$ and $E=\{e_{1},e_{2},\ldots,e_{m}\}$. For each vertex $v_{i}\in V$, we consider two $tcmbt_{3}$ with $v_{i}$ and $v_{i}^{\prime}$ as their roots, respectively. Similarly, for each edge $e_{i}\in E$, we consider another two $tcmbt_{3}$ with $v_{e_{i}}$ and $v_{e_{i}}^{\prime}$ as their roots, respectively. Further, corresponding to every edge $e_{j}\in E$, we take two vertices $e_{j},e_{j}^{\prime}$ in $G^{\prime}$. Also, we take another two vertices $e$ and $e^{\prime}$ in $G^{\prime}$. Let $A=\{e_{1},e_{2},\ldots,e_{m},e\}$ and $B=\{e_{1}^{\prime},e_{2}^{\prime},\ldots,e_{m}^{\prime},e^{\prime}\}$. We construct a complete bipartite graph with vertex set $A\cup B$. Now, for each edge $e_{t}=v_{i}v_{j}\in E$, we join the edges $v_{i}e_{t}$, $v_{j}e_{t}$, $v_{e_{t}}e_{t}$, $v_{i}^{\prime}e_{t}^{\prime}$, $v_{j}^{\prime}e_{t}^{\prime}$, and $v_{e_{t}}^{\prime}e_{t}^{\prime}$ in $G^{\prime}$. Next, we consider six $tcmbt_{3}$ with roots $v_{a},v_{a}^{\prime},v_{e},v_{e}^{\prime},v_{b},v_{b}^{\prime}$. Finally, we make $\{v_{a},v_{e},v_{b}\}$, $\{v_{a}^{\prime},v_{e}^{\prime},v_{b}^{\prime}\}$ adjacent to the vertex $e$, and $e^{\prime}$, respectively, and we set $k=m+4$. In this construction, we can verify that the graph $G^{\prime}=(V^{\prime},E^{\prime})$ consists of $11640m+38n+$ 116 vertices and $m^{2}+42m+34n+104$ edges. The construction is illustrated in Figure3.
@@
Let $g$ be a proper $3$-coloring of $G$. Based on $g$, let us consider a vertex partition, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, of $G^{\prime}$ as follows: for each $v_{i}\in V$, we put the corresponding vertex $v_{i},v_{i}^{\prime}$ in $V_{q}$, if $g(v_{i})=q$. For each edge $e_{t}=v_{i}v_{j}\in E$, we put the vertex $v_{e_{t}},v_{e_{t}}^{\prime}\in V_{q}$, if $g(v_{i})\neq q$ and $g(v_{j})\neq q$. Further, we put $v_{a},v_{a}^{\prime}\in V_{3}$, $v_{e},v_{e}^{\prime}\in V_{2}$, and $v_{b},v_{b}^{\prime}\in V_{1}$. We put the other vertices of $tcmbt_{3}$ according to Figure 4, based on the position of the root vertex in the partition. Finally, for $1\leq j\leq m$, we put $e_{j},e_{j}^{\prime}\in V_{3+j}$ and $e,e^{\prime}\in V_{m+4}$.
Figure 4: Partition of $tcmbt_{3}$
Since the vertices of $A\cup B$ induces a complete bipartite graph, $V_{i}$ dominates $V_{j}$ for $4\leq i\leq j\leq k$. Also, for each $i=1,2,3$, every vertex of $A\cup B$ is adjacent to two vertices of $V_{i}$. Therefore, for each $i=1,2,3$, $V_{i}$ dominates $V_{t}$, for all $t>3$. Finally, from Figure4, it is evident that $V_{i}$ dominates $V_{j}$ for $1\leq i\leq j\leq 3$. Hence, $\pi$ is a total transitive partition of $G^{\prime}$ of size $k$. Therefore, if $G$ has a proper $3$-coloring, then $G^{\prime}$ has a total transitive partition of size $k$. ∎
@@
First note that the vertices $\{v_{a},v_{e},v_{b}\}$, which are adjacent to $e$, must belong to $V_{1},V_{2}$ or $V_{3}$. As the degree of each vertex from $\{v_{a},v_{e},v_{b}\}$ is $4$, they can be at most in $V_{4}$, but in this case, $e$ can be either in $V_{3}$ or $V_{4}$. Furthermore, if $e$ is in $V_{3}$, there is no neighbour in $V_{4}$ to dominate vertices from $\{v_{a},v_{e},v_{b}\}$ and similarly if $e$ is in $V_{4}$, then here is no neighbour in $V_{3}$ to dominate vertices from $\{v_{a},v_{e},v_{b}\}$. Similarly, the vertices of $\{v^{\prime}_{a},v^{\prime}_{e},v^{\prime}_{b}\}$, which are adjacent to $e^{\prime}$ cannot be in $V_{p}$ for any $p\geq 4$. Further, since the degree of every vertex other than the vertices of $A\cup B$ is less than $m+4$, $V_{k}$ can contain two vertices only from $A\cup B=\{e_{1},e_{2},\ldots,e_{m},e,e_{1}^{\prime},e_{2}^{\prime},\ldots,e_{m}^{\prime},e^{\prime}\}$. In this proof, we show that to dominate the vertices of $A\cup B$ which belong to $V_{k}$, the other vertices of $A\cup B$ must be in $V_{4},V_{5},\ldots,V_{k-1}$. This, in turn, would imply that the sets $V_{1},V_{2}$, and $V_{3}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$. The proof is divided into four cases based on whether $e\in V_{k}$ or not and $e^{\prime}\in V_{k}$ or not.
@@
As discussed earlier, the vertices $\{v_{a},v_{e},v_{b}\}$ and $\{v^{\prime}_{a},v^{\prime}_{e},v^{\prime}_{b}\}$ belong to $V_{p}$ for some $1\leq p\leq 3$. Since $e\in V_{k}$ to dominate $e$, each set in $\{V_{1},\ldots,V_{k-1},V_{k}\}$ must contain at least one vertices from $N_{G^{\prime}}(e)=\{e_{1}^{\prime},e_{2}^{\prime},\ldots,e_{m}^{\prime},e^{\prime},v_{a},v_{e},v_{b}\}$. Since $e$ is adjacent to exactly $k=m+4$ vertices, each $V_{i},1\leq i\leq k$ contains exactly one vertex from $N_{G^{\prime}}(e)$. Since $\{v_{a},v_{e},v_{b}\}$ belong to $V_{p}$ for some $1\leq p\leq 3$ it follows that vertices from $N_{G^{\prime}}(e)\setminus\{v_{a},v_{e},v_{b}\}$ belong to $V_{4},V_{5},\ldots V_{k}$. Similarly, as $e^{\prime}\in V_{k}$, to dominate $e^{\prime}$, each set in $\{V_{1},V_{2},\ldots,V_{k-1},V_{k}\}$ must contain at least one vertex from $N_{G^{\prime}}(e^{\prime})=\{e_{1},e_{2},\ldots,e,v_{a}^{\prime},v_{e}^{\prime},v_{b}^{\prime}\}$. Since $e^{\prime}$ is adjacent to exactly $k=m+4$ vertices, each $V_{i},1\leq i\leq k$ contains exactly one vertex from $N_{G^{\prime}}(e^{\prime})$. Similarly, as before, since $\{v_{a}^{\prime},v_{e}^{\prime},v_{b}^{\prime}\}$ belong to $V_{p}$ for some $1\leq p\leq 3$, it follows that exactly three vertices from $N_{G^{\prime}}(e^{\prime})\setminus\{v^{\prime}_{a},v^{\prime}_{e},v^{\prime}_{b}\}$ belong to $V_{4},V_{5},\ldots V_{k}$. Hence, the vertices of $A\cup B$ belong to $\{V_{4},V_{5},\ldots V_{k}\}$.
@@
Using similar arguments as in the previous case, we know that vertices from $\{v_{a},v_{b},v_{e}\}$ belong to $V_{p}$ for $1\leq p\leq 3$ and the vertices from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m},e^{\prime}\}$ belong to $\{V_{4},V_{5},\ldots,V_{k-1},V_{k}\}$ to dominate $e$. Moreover, every $V_{i}$ for $4\leq i\leq k$ contains exactly one vertex from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m},e^{\prime}\}$. Since the vertices from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m},e^{\prime}\}$ belong to $\{V_{4},V_{5},\ldots,V_{k-1},V_{k}\}$, no vertex from $\{v^{\prime}_{1},\ldots,v^{\prime}_{n},v^{\prime}_{e_{1}},\ldots,v^{\prime}_{e_{m}},v^{\prime}_{a},v^{\prime}_{b},v^{\prime}_{e}\}$ can be in $V_{p}$ for $p\geq 4$. As $e^{\prime}\notin V_{k}$, there exist a vertex from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m}\}$, say $e^{\prime}_{r}$ such that $e^{\prime}_{r}\in V_{k}$. To dominate $e^{\prime}_{r}$, the vertices from $\{e_{1},e_{2},\ldots,e_{m},e\}$ belong to $\{V_{4},V_{5},\ldots,V_{k}\}$. With similar arguments as in the previous case, we know that every vertex of $\{v_{1},v_{2},\ldots,v_{n},v_{e_{1}},v_{e_{2}},\ldots,v_{e_{m}}\}$ belong to $V_{p}$ for $1\leq p\leq 3$. Hence, $\{V_{4},V_{5},\ldots V_{k}\}$ contain only vertices from $A\cup B$ and $\{V_{1},V_{2},V_{3}\}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$.
@@
Since the degree of each vertex of $V^{\prime}\setminus(A\cup B)$ is at most $k-1=m+3$, they cannot belong to $V_{k}$, that is, only vertices from $A\cup B$ can be in $V_{k}$. Without loss of generality, we can assume that $e_{1}\in V_{k}$ and $e^{\prime}_{s}\in V_{k}$ in the total transitive partition of $G^{\prime}$ of size $k$. Also, let $e_{1}$ be the edge between $v_{1}$ and $v_{2}$ in $G$ and in the total transitive partition of $G^{\prime}$ of size $k$, $v_{1}\in V_{l}$ and $v_{2}\in V_{t}$ where $l\geq t$. Next, we show that $l\leq 3$. As $v_{1}\in V_{l}$ and $\pi$ is a total transitive partition of $G$ implies that each of $V_{3},V_{4},\ldots,V_{l}$ contains at least one vertex from $A\setminus\{e_{1},e\}$. On the other hand, as $e_{s}^{\prime}\in V_{k}$, each of $\{V_{l+1},\ldots,V_{k-1}\}$ also contains at least one vertex from $A\setminus\{e_{1}\}$. This implies that each of the $(m+1)$ partitions, $\{V_{3},V_{4},\ldots,V_{k-1}\}$, contains one vertex from the set of $m$ vertices, $A\setminus\{e_{1}\}$, which is impossible. Hence, $l\leq 3$. With a similar argument, we can show that $v_{e_{1}}$ cannot be in $V_{p}$ for $p\geq 4$. Therefore, each $\{V_{3},V_{4},\ldots,V_{k}\}$ contains exactly one vertex from $B$. Since the vertices of $B$ belong to $\{V_{4},V_{5},\ldots,V_{k}\}$, the vertices of $\{v^{\prime}_{1},v^{\prime}_{2},\ldots,v^{\prime}_{n},v^{\prime}_{e_{1}},v^{\prime}_{e_{2}},\ldots,v^{\prime}_{e_{m}},v^{\prime}_{a},v^{\prime}_{b},v^{\prime}_{e}\}$ belong to some partition $V_{p}$ for some $p\leq 3$. As $e^{\prime}_{s}\in V_{k}$ and $\pi$ is a total transitive partition of $G$ to dominate $e^{\prime}_{s}$, the vertices from $A\setminus\{e_{1}\}$ belong to $\{V_{4},V_{5},\ldots,V_{k-1}\}$. With similar arguments as in Case 1, we know that each vertex of $\{v_{1},v_{2},\ldots,v_{n},v_{e_{1}},v_{e_{2}},\ldots,v_{e_{m}},v_{a},v_{b},v_{e}\}$ belongs to $V_{p}$ for some $p\leq 3$. Hence, $\{V_{4},V_{5},\ldots V_{k}\}$ contain only vertices from $A\cup B$ and $\{V_{1},V_{2},V_{3}\}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$.

### v2

@@ 3 Total Transitivity in split graphs
$|dom_{K}(S)|\geq 2$
In this case, let $l_{min}=|dom_{K}(S)|$. By the definition of $dom_{K}(S)$, the vertices are in $V_{1}$. As we know from the Observation 14, each $V_{i}$ must contain a vertex from $K$, so each $(p-1)-[q-\{(p-1)+l_{min}\}]=2p-q+l_{min}-2$ set from $\pi$ must contain at least one vertex from $S$. Therefore, $|S|\geq 2p-q+l_{min}-2$ $(b(i))$. Again, by Observation 15, each $|V_{i}\cap S|\leq 2$. Let $\{s_{i_{1}},s_{i_{2}},\ldots,s_{i_{\beta}}\}\subseteq S$, where $\beta=2p-q+l_{min}-2$ such that $s_{i_{j}}\in V_{i_{j}}$. Also, note that $i_{j}\geq 2$. Since $\pi$ is a total transitive partition, $deg(s_{i_{j}})\geq i_{j}\geq j$ $(b(ii))$.
@@
Next, we show that the above conditions are not sufficient for $Tr_{t}(G)=p$. Let us see an example of a graph given in the Figure 3.
Figure 3: Example of a split graph
In the above graph $G$, $q=\omega(G)=8$. We demonstrate that for $p=6$, $G$ satisfies condition $(a)$ of Theorem 19, but $Tr_{t}(G)\neq 6$. Clearly, a $dom_{K}(S)$-set exists, and $dom_{K}(S)=\{x_{1}\}$, which implies $|dom_{K}(S)|=1$. Also, $84=|S|\geq 2p-q=12-$ 8. Moreover, $\alpha=2p-q=4$, and $\{s_{1},s_{2},s_{3},s_{4}\}\subseteq S$ such that $\deg(s_{i})\geq i$. Thus, $G$ satisfies condition $(a)$ of Theorem 19.
Next, we show that $Tr_{t}(G)\neq 6$. Assume, for contradiction, that $Tr_{t}(G)=6$, and let $\pi=\{V_{1},V_{2},V_{3},V_{4},V_{5},V_{6}\}$ be a total transitive partition of $G$ of size $6$. By the definition of $dom_{K}(S)$, we know that $x_{1}\in V_{1}$. From the degree of $s_{i}$, it follows that $s_{i}\notin V_{j}$ for $j\geq 5$. Since $\pi$ is a total transitive partition, $|V_{i}|\geq 2$, and by Observation 14, each $V_{i}$ must contain a vertex from $K$. Therefore, each $V_{i}$ contains exactly one vertex from $K=\{x_{1},x_{2},\ldots,x_{7},x_{8}\}$ for $1\leq i\leq 4$. Additionally, from the degrees of $s_{i}$, we deduce that $s_{i}\in V_{i}$ for all $1\leq i\leq 4$. To totally dominate $s_{4}$, we have $\{x_{1},x_{2},x_{3},x_{5}\}\subseteq\bigcup_{i=1}^{4}V_{i}$, and to totally dominate $s_{3}$, $\{x_{1},x_{6},x_{7}\}\subseteq\bigcup_{i=1}^{3}V_{i}$. This leads to a contradiction, as $|K\cap\bigcup_{i=1}^{4}V_{i}|=4$.
Therefore, the total transitivity of $G$ cannot be $6$. Hence, condition $(a)$ of Theorem 19 is not sufficient for $Tr_{t}(G)=p$.
@@ 4 NP-completeness
In this section, we show that the Maximum Total Transitivity Decision Problem is NP-complete for a bipartite graph. Clearly, this problem is in NP. We prove the NP-hardness of this problem by showing a polynomial-time reduction from Proper $3$-Coloring Decision Problem in graphs to the Maximum Total Transitivity Decision Problem. A proper $3$-colring of a graph $G=(V,E)$ is a function $g:V\rightarrow\{1,2,3\}$, such that for every edge $uv\in E$, $g(u)\not=g(v)$. The proper $3$-Coloring Decision Problem is defined as follows:
@@
Figure 4: Construction of $G^{\prime}$ from $G$ where $m=3$ and $T$ represents a $\mathcal{T}_{3}$
The construction is as follows. Let $V=\{v_{1},v_{2},\ldots,v_{n}\}$ and $E=\{e_{1},e_{2},\ldots,e_{m}\}$. Let us consider the total transitive tree of order $3$, that is $\mathcal{T}_{3}$. For each vertex $v_{i}\in V$, we consider two $\mathcal{T}_{3}$ with $v_{i}$ and $v_{i}^{\prime}$ as their roots, respectively. Similarly, for each edge $e_{i}\in E$, we consider another two $\mathcal{T}_{3}$ with $v_{e_{i}}$ and $v_{e_{i}}^{\prime}$ as their roots, respectively. Further, corresponding to every edge $e_{j}\in E$, we take two vertices $e_{j},e_{j}^{\prime}$ in $G^{\prime}$. Also, we take another two vertices $e$ and $e^{\prime}$ in $G^{\prime}$. Let $A=\{e_{1},e_{2},\ldots,e_{m},e\}$ and $B=\{e_{1}^{\prime},e_{2}^{\prime},\ldots,e_{m}^{\prime},e^{\prime}\}$. We construct a complete bipartite graph with vertex set $A\cup B$. Now, for each edge $e_{t}=v_{i}v_{j}\in E$, we join the edges $v_{i}e_{t}$, $v_{j}e_{t}$, $v_{e_{t}}e_{t}$, $v_{i}^{\prime}e_{t}^{\prime}$, $v_{j}^{\prime}e_{t}^{\prime}$, and $v_{e_{t}}^{\prime}e_{t}^{\prime}$ in $G^{\prime}$. Next, we consider six $\mathcal{T}_{3}$ with roots $v_{a},v_{a}^{\prime},v_{e},v_{e}^{\prime},v_{b},v_{b}^{\prime}$. Finally, we make $\{v_{a},v_{e},v_{b}\}$, $\{v_{a}^{\prime},v_{e}^{\prime},v_{b}^{\prime}\}$ adjacent to the vertex $e$, and $e^{\prime}$, respectively, and we set $k=m+4$. In this construction, we can verify that the graph $G^{\prime}=(V^{\prime},E^{\prime})$ consists of $11640m+38n+$ 116 vertices and $m^{2}+42m+34n+104$ edges. The construction is illustrated in Figure 4.
@@
Let $g$ be a proper $3$-coloring of $G$. Based on $g$, let us consider a vertex partition, say $\pi=\{V_{1},V_{2},\ldots,V_{k}\}$, of $G^{\prime}$ as follows: for each $v_{i}\in V$, we put the corresponding vertex $v_{i},v_{i}^{\prime}$ in $V_{q}$, if $g(v_{i})=q$. For each edge $e_{t}=v_{i}v_{j}\in E$, we put the vertex $v_{e_{t}},v_{e_{t}}^{\prime}\in V_{q}$, if $g(v_{i})\neq q$ and $g(v_{j})\neq q$. Further, we put $v_{a},v_{a}^{\prime}\in V_{3}$, $v_{e},v_{e}^{\prime}\in V_{2}$, and $v_{b},v_{b}^{\prime}\in V_{1}$. We put the other vertices of $\mathcal{T}_{3}$ according to Figure 5, based on the position of the root vertex in the partition. Finally, for $1\leq j\leq m$, we put $e_{j},e_{j}^{\prime}\in V_{3+j}$ and $e,e^{\prime}\in V_{m+4}$.
Figure 5: Partition of $\mathcal{T}_{3}$
Since the vertices of $A\cup B$ induces a complete bipartite graph, $V_{i}$ totally dominates $V_{j}$ for $4\leq i\leq j\leq k$. Also, for each $i=1,2,3$, every vertex of $A\cup B$ is adjacent to two vertices of $V_{i}$. Therefore, for each $i=1,2,3$, $V_{i}$ totally dominates $V_{t}$, for all $t>3$. Finally, from Figure 5, it is evident that $V_{i}$ totally dominates $V_{j}$ for $1\leq i\leq j\leq 3$. Hence, $\pi$ is a total transitive partition of $G^{\prime}$ of size $k$. Therefore, if $G$ has a proper $3$-coloring, then $G^{\prime}$ has a total transitive partition of size $k$. ∎
@@
First note that the vertices $\{v_{a},v_{e},v_{b}\}$, which are adjacent to $e$, must belong to $V_{1},V_{2}$ or $V_{3}$. As the degree of each vertex from $\{v_{a},v_{e},v_{b}\}$ is $4$, they can be at most in $V_{4}$, but in this case, $e$ can be either in $V_{3}$ or $V_{4}$. Furthermore, if $e$ is in $V_{3}$, there is no neighbour in $V_{4}$ to totally dominate vertices from $\{v_{a},v_{e},v_{b}\}$ and similarly if $e$ is in $V_{4}$, then here is no neighbour in $V_{3}$ to totally dominate vertices from $\{v_{a},v_{e},v_{b}\}$. Similarly, the vertices of $\{v^{\prime}_{a},v^{\prime}_{e},v^{\prime}_{b}\}$, which are adjacent to $e^{\prime}$ cannot be in $V_{p}$ for any $p\geq 4$. Further, since the degree of every vertex other than the vertices of $A\cup B$ is less than $m+4$, $V_{k}$ can contain two vertices only from $A\cup B=\{e_{1},e_{2},\ldots,e_{m},e,e_{1}^{\prime},e_{2}^{\prime},\ldots,e_{m}^{\prime},e^{\prime}\}$. In this proof, we show that to totally dominate the vertices of $A\cup B$ which belong to $V_{k}$, the other vertices of $A\cup B$ must be in $V_{4},V_{5},\ldots,V_{k-1}$. This, in turn, would imply that the sets $V_{1},V_{2}$, and $V_{3}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$. The proof is divided into four cases based on whether $e\in V_{k}$ or not and $e^{\prime}\in V_{k}$ or not.
@@
As discussed earlier, the vertices $\{v_{a},v_{e},v_{b}\}$ and $\{v^{\prime}_{a},v^{\prime}_{e},v^{\prime}_{b}\}$ belong to $V_{p}$ for some $1\leq p\leq 3$. Since $e\in V_{k}$ to totally dominate $e$, each set in $\{V_{1},\ldots,V_{k-1},V_{k}\}$ must contain at least one vertices from $N_{G^{\prime}}(e)=\{e_{1}^{\prime},e_{2}^{\prime},\ldots,e_{m}^{\prime},e^{\prime},v_{a},v_{e},v_{b}\}$. Since $e$ is adjacent to exactly $k=m+4$ vertices, each $V_{i},1\leq i\leq k$ contains exactly one vertex from $N_{G^{\prime}}(e)$. Since $\{v_{a},v_{e},v_{b}\}$ belong to $V_{p}$ for some $1\leq p\leq 3$ it follows that vertices from $N_{G^{\prime}}(e)\setminus\{v_{a},v_{e},v_{b}\}$ belong to $V_{4},V_{5},\ldots V_{k}$. Similarly, as $e^{\prime}\in V_{k}$, to totally dominate $e^{\prime}$, each set in $\{V_{1},V_{2},\ldots,V_{k-1},V_{k}\}$ must contain at least one vertex from $N_{G^{\prime}}(e^{\prime})=\{e_{1},e_{2},\ldots,e,v_{a}^{\prime},v_{e}^{\prime},v_{b}^{\prime}\}$. Since $e^{\prime}$ is adjacent to exactly $k=m+4$ vertices, each $V_{i},1\leq i\leq k$ contains exactly one vertex from $N_{G^{\prime}}(e^{\prime})$. Similarly, as before, since $\{v_{a}^{\prime},v_{e}^{\prime},v_{b}^{\prime}\}$ belong to $V_{p}$ for some $1\leq p\leq 3$, it follows that exactly three vertices from $N_{G^{\prime}}(e^{\prime})\setminus\{v^{\prime}_{a},v^{\prime}_{e},v^{\prime}_{b}\}$ belong to $V_{4},V_{5},\ldots V_{k}$. Hence, the vertices of $A\cup B$ belong to $\{V_{4},V_{5},\ldots V_{k}\}$.
@@
Using similar arguments as in the previous case, we know that vertices from $\{v_{a},v_{b},v_{e}\}$ belong to $V_{p}$ for $1\leq p\leq 3$ and the vertices from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m},e^{\prime}\}$ belong to $\{V_{4},V_{5},\ldots,V_{k-1},V_{k}\}$ to totally dominate $e$. Moreover, every $V_{i}$ for $4\leq i\leq k$ contains exactly one vertex from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m},e^{\prime}\}$. Since the vertices from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m},e^{\prime}\}$ belong to $\{V_{4},V_{5},\ldots,V_{k-1},V_{k}\}$, no vertex from $\{v^{\prime}_{1},\ldots,v^{\prime}_{n},v^{\prime}_{e_{1}},\ldots,v^{\prime}_{e_{m}},v^{\prime}_{a},v^{\prime}_{b},v^{\prime}_{e}\}$ can be in $V_{p}$ for $p\geq 4$. As $e^{\prime}\notin V_{k}$, there exist a vertex from $\{e^{\prime}_{1},e^{\prime}_{2},\ldots,e^{\prime}_{m}\}$, say $e^{\prime}_{r}$ such that $e^{\prime}_{r}\in V_{k}$. To totally dominate $e^{\prime}_{r}$, the vertices from $\{e_{1},e_{2},\ldots,e_{m},e\}$ belong to $\{V_{4},V_{5},\ldots,V_{k}\}$. With similar arguments as in the previous case, we know that every vertex of $\{v_{1},v_{2},\ldots,v_{n},v_{e_{1}},v_{e_{2}},\ldots,v_{e_{m}}\}$ belong to $V_{p}$ for $1\leq p\leq 3$. Hence, $\{V_{4},V_{5},\ldots V_{k}\}$ contain only vertices from $A\cup B$ and $\{V_{1},V_{2},V_{3}\}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$.
@@
Since the degree of each vertex of $V^{\prime}\setminus(A\cup B)$ is at most $k-1=m+3$, they cannot belong to $V_{k}$, that is, only vertices from $A\cup B$ can be in $V_{k}$. Without loss of generality, we can assume that $e_{1}\in V_{k}$ and $e^{\prime}_{s}\in V_{k}$ in the total transitive partition of $G^{\prime}$ of size $k$. Also, let $e_{1}$ be the edge between $v_{1}$ and $v_{2}$ in $G$ and in the total transitive partition of $G^{\prime}$ of size $k$, $v_{1}\in V_{l}$ and $v_{2}\in V_{t}$ where $l\geq t$. Next, we show that $l\leq 3$. As $v_{1}\in V_{l}$ and $\pi$ is a total transitive partition of $G$ implies that each of $V_{3},V_{4},\ldots,V_{l}$ contains at least one vertex from $A\setminus\{e_{1},e\}$. On the other hand, as $e_{s}^{\prime}\in V_{k}$, each of $\{V_{l+1},\ldots,V_{k-1}\}$ also contains at least one vertex from $A\setminus\{e_{1}\}$. This implies that each of the $(m+1)$ partitions, $\{V_{3},V_{4},\ldots,V_{k-1}\}$, contains one vertex from the set of $m$ vertices, $A\setminus\{e_{1}\}$, which is impossible. Hence, $l\leq 3$. With a similar argument, we can show that $v_{e_{1}}$ cannot be in $V_{p}$ for $p\geq 4$. Therefore, each $\{V_{3},V_{4},\ldots,V_{k}\}$ contains exactly one vertex from $B$. Since the vertices of $B$ belong to $\{V_{4},V_{5},\ldots,V_{k}\}$, the vertices of $\{v^{\prime}_{1},v^{\prime}_{2},\ldots,v^{\prime}_{n},v^{\prime}_{e_{1}},v^{\prime}_{e_{2}},\ldots,v^{\prime}_{e_{m}},v^{\prime}_{a},v^{\prime}_{b},v^{\prime}_{e}\}$ belong to some partition $V_{p}$ for some $p\leq 3$. As $e^{\prime}_{s}\in V_{k}$ and $\pi$ is a total transitive partition of $G$ to totally dominate $e^{\prime}_{s}$, the vertices from $A\setminus\{e_{1}\}$ belong to $\{V_{4},V_{5},\ldots,V_{k-1}\}$. With similar arguments as in Case 1, we know that each vertex of $\{v_{1},v_{2},\ldots,v_{n},v_{e_{1}},v_{e_{2}},\ldots,v_{e_{m}},v_{a},v_{b},v_{e}\}$ belongs to $V_{p}$ for some $p\leq 3$. Hence, $\{V_{4},V_{5},\ldots V_{k}\}$ contain only vertices from $A\cup B$ and $\{V_{1},V_{2},V_{3}\}$ contain only vertices from $V^{\prime}\setminus(A\cup B)$.